OpenStudy (emilyisanemoji):

Will medal best answer and possibly fan! what property is this? (34 + 11) + 9 = (11 + 34) + 9 a. commutative----\ b. associative------ >properties c. opposite of a sum/

OpenStudy (abbles):

Associative property

OpenStudy (emilyisanemoji):

what about 25 + (-15 + 4) = (25 + -15)) + 4

OpenStudy (abbles):

That would be the Associative Property as well.

OpenStudy (abbles):

When adding/subtracting, you are just switching the numbers that the parenthesis are around. In the Order of Operations, this is allowed in addition/subtraction but not multiplication or division.

OpenStudy (abbles):

Do you understand?

OpenStudy (emilyisanemoji):

and -(-21 + 8) = 21 + (-8)

OpenStudy (abbles):

Take a look at this http://www.purplemath.com/modules/numbprop.htm

OpenStudy (mjdennis):

Um, @Abbles is incorrect. Look closely and you will see your problem is not the same as his/her example. "Associate" and "associative" means "who you hang out with". Your problem never changes what pnumbers are grouped together, it just changes the order, so it is not associative

OpenStudy (mjdennis):

@Abbles is, however, correct on the second one.

OpenStudy (emilyisanemoji):

then what is it on the first?

OpenStudy (abbles):

@mjdennis The first equation is the Associative Property. ?

OpenStudy (abbles):

Associative Property: a + (b + c) = (a + b) + c

OpenStudy (abbles):

Commutative Property: a + b = b + a

jimthompson5910 (jim_thompson5910):

I agree with @mjdennis `(34 + 11) + 9 = (11 + 34) + 9` is commutative. The terms 11 and 34 swap places and change the order of addition `25 + (-15 + 4) = (25 + -15) + 4` is associative. The parenthesis move around to form another group.

OpenStudy (abbles):

Ah I see. Didn't look closely enough at the numbers. So it's Commutative because the numbers in the parenthesis don't change?

OpenStudy (mjdennis):

@Abbles , like I said, the original problem does not move the groupings around, it is NOT associative. It moves the order inside the parentheses. That is a different property. @Emilyisanemoji , if I wanted to ell you outright, I would have done so. I wanted you to go back and read the stuff that @Abbles was trying to use to teach you, and let you think for yourself.

OpenStudy (emilyisanemoji):

what about -(-21 + 8) = 21 + (-8)

OpenStudy (mjdennis):

Pretend that it is (-1) * (-21 + 8) instead. How did they move the -1 to both parts of the answer = 21 + (-8) ?

OpenStudy (emilyisanemoji):

umm...............they multiply?.....

OpenStudy (mjdennis):

@Abbles , yep, commutative. _Commuters_ move to and from work. Here in math, commutative means we don't care where we move things having the same operation, like addition

OpenStudy (emilyisanemoji):

im not following...

OpenStudy (mjdennis):

@Emilyisanemoji multiply is part of it. On the left, there is only one -1. On the right, the -1 has been spread over two numbers to multiply them. It's a tough one, but the property is another meaning for "to spread out"

OpenStudy (mjdennis):

To ... distribute? Distribut-ive property?

OpenStudy (emilyisanemoji):

im sorry, but i really, truley dont understand. distributive property isnt one of my answers.

OpenStudy (mjdennis):

OH. Oops. Let me think

OpenStudy (mjdennis):

Google says "The opposite_of_a_sum of real numbers is equal to the _sum_ of the _opposites_." (hotmath.com/hotmath_help/topics/properties-of-addition.html) -(-21 + 8) is the opposite of (-21 + 8) 21 is the opposite of -2, and +8 is the opposite of -8

OpenStudy (mjdennis):

I mean, 21 is the opposite of -21

OpenStudy (emilyisanemoji):

and.....?

OpenStudy (emilyisanemoji):

hello?

OpenStudy (mjdennis):

And the fact that I used the words "the opposite_of_a_sum" really gives you no clue as to which of your choices is correct?

OpenStudy (mjdennis):

Go back and read your choices.

OpenStudy (emilyisanemoji):

all it is